OpenStudy (anonymous):

What are all the factors of 32?

OpenStudy (geometry_hater):

32 16 12 9 8 6 4 3 2 1

OpenStudy (anonymous):

umm 9 isn't a factor

OpenStudy (geometry_hater):

oops my bad

OpenStudy (anonymous):

and 12 isn't a factor too :-P

OpenStudy (anonymous):

3 isn't a factor :-P

OpenStudy (anonymous):

but all the rest are indeed factors of 32

OpenStudy (anonymous):

Make a factor tree (I tried my best to format one below). \[32\left\{\begin{array}{l}2 \\ 16\left\{\begin{array}{l}2 \\ 8\left\{\begin{array}{l}2 \\ 4\left\{\begin{array}{l}2 \\ 2\end{array}\right.\end{array}\right.\end{array}\right.\end{array}\right.\]Any combinations of these prime factors will be factors of 32. Since they're all 2s, the process is easy: \[\left\{1,2,2^2,2^3,2^4,2^5\right\}=\boxed{\left\{1, 2, 4, 8, 16, 32\right\}}\]

OpenStudy (anonymous):

except for 6